OpenStudy (anonymous):

PLZ HELP WITH THIS ALGEBRA IVE BEEN WAITING FOREVER FOR SOMEONE TO HELP ME! equation is below

OpenStudy (anonymous):

\[ \frac{ 4 \sqrt{150} }{ \sqrt{189}}\]

OpenStudy (anonymous):

189x*

OpenStudy (anonymous):

not 189